OpenStudy (anonymous):

Subtract and simplify:

OpenStudy (anonymous):

\[\frac{ x+3 }{ 5 }-\frac{ 2x+1 }{ 10 }\]

OpenStudy (anonymous):

we have to take l.c.m first, 2(x+3)-(2x+1)/10 =2x+6-2x-1/10 =5/10 =1/2

OpenStudy (anonymous):

have u understood each and every step?ask me if u hv any query

OpenStudy (anonymous):

thank you :) im sure if I get it all thou lol